solve..6(x+4)=8(x-2)...x=???
Answer by Jstrasner(112)

Hey,
So for this one you need to distribute the numbers outside of the parentheses:
6(x+4)=8(x-2) => 6x + 24 = 8x - 16
Now we subtract 6x from both sides and add 16 to both sides:
6x - 6x + 24 + 16 = 8x - 6x - 16 + 16 =>
40 = 2x
then we divide both sides by 2:
x = 20
I hope this helps!